OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s^2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal.

OpenStudy (anonymous):

also to the nearest whole number as well. I'm so confused. help!

OpenStudy (anonymous):

I this question in your parabola or quadratic equation chapter

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

both

OpenStudy (anonymous):

have you bee using graphing calculators

OpenStudy (anonymous):

use you graphing calc and plug in for y= your equation

OpenStudy (anonymous):

dont have one.

OpenStudy (anonymous):

hmmm...oh, you just need to max value for s, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

remember, the graph will point downwards since the leading term has a -.....remember that?

OpenStudy (anonymous):

so you need the max s value which at the vertex...top of the parabola when sketched out....lkooks like the path when you throw a ball up and it comes back down...got the picture?

OpenStudy (anonymous):

so you can find the x value of the vertex by using -a/2b....where a is the coeficient of the leading term and b is the coefficient of the second, or linear, term.

OpenStudy (anonymous):

ok make since so far.

OpenStudy (anonymous):

so you'll have -(-0.025)/2(2.45)......have a regular calculator?

OpenStudy (anonymous):

that's backwards.....-b/2a = -2.45/2(-0.025)= 49 so plug in the 49 for s and evaluate....

OpenStudy (anonymous):

are you still there??